OpenStudy (anonymous):

Simplify f(x) =(2x^2+3x+1)/(2x+1). Using complete sentences, explain why f(1) = 2, f(0) = 1, and f(-1) = 0, yet f(-1/2) is undefined. Be sure to show your work.

OpenStudy (anonymous):

(2x^2+3x+1)/(2x+1). (2x+1) (x + 1)/(2x+1) (x+1)

OpenStudy (anonymous):

So, f(x) = (x+1) if there is f(1) then replace x by 1 and further

OpenStudy (anonymous):

for..., let see explain why f(1) = 2,??? because f(x) = x + 1 f(1) = 1 + 1 f(1) = 2